1) A company is considering the purchase of new equipment for $90,000. The projected annual net...

1)

A company is considering the purchase of new equipment for $90,000. The projected annual net cash flows are $35,500. The machine has a useful life of 3 years and no salvage value. Management of the company requires a 8% return on investment. The present value of an annuity of $1 for various periods follows:
    

Period Present value of an annuity of $1 at 8%
1 0.9259
2 1.7833
3 2.5771

   
What is the net present value of this machine assuming all cash flows occur at year-end?

Multiple Choice

  • $30,000

  • $4,500

  • $1,487

  • $34,500

  • $88,910

2)

The following present value factors are provided for use in this problem.

Periods Present Value
of $1 at 8%
Present Value of an
Annuity of $1 at 8%
1 0.9259 0.9259
2 0.8573 1.7833
3 0.7938 2.5771
4 0.7350 3.3121


Xavier Co. wants to purchase a machine for $36,900 with a four year life and a $1,200 salvage value. Xavier requires an 8% return on investment. The expected year-end net cash flows are $11,900 in each of the four years. What is the machine's net present value?

Multiple Choice

  • $3,396.

  • $2,514.

  • $40,296.

  • $(3,396).

  • $(2,514).

3)

A company is considering the purchase of a new machine for $57,000. Management predicts that the machine can produce sales of $16,900 each year for the next 10 years. Expenses are expected to include direct materials, direct labor, and factory overhead totaling $7,100 per year including depreciation of $4,900 per year. Income tax expense is $3,920 per year based on a tax rate of 40%. What is the payback period for the new machine?

Multiple Choice

  • 3.37 years.

  • 6.40 years.

  • 5.29 years.

  • 11.63 years.
Answer #1

1.

Net present value = Present value of cash inflow - Initial investment

Net present value = $35,500*2.5771 - $90,000

Net present value = $91,487 - 90,000 = $1,487

2.

Net present value = $11,900*3.3121 + 1,200*0.7350 - $36,900

Net present value = $39,414 + 882 - $36,900 = $3,396

3.

Pay back period = Initial investment / Annual cash inflow

Annual cash inflow = $16,900 - 7,100 - 3,920 + 4,900 = $10,780

Pay back period = $57,000 / 10,780 = 5.29 years